Mirror Cover and Dome Rotation Failures

Tuesday, November 15, 2016 10:32 AM

Problem(s) Encountered:

As mentioned the mirror covers closed in the wrong order W E on top N S

Called Eric and helped him pull covers open.  

After Eric left went to first target and dome continued to spin CW.  Tried a couple of solutions but finally sent dome home, cycled power and reset @ home to 320 that worked for a few targets. This happened several times throughout the night, resetting seemed to fix the issue for a few targets then back to spinning.  When sending the dome to home it was off by 20.

Stopped trying to fix dome and just went to moving dome manually for rest of night.


Since both, the east and west petals had closed incorrectly, there was simply no way for a single person to resolve the issue.  Indeed even with two people, it is a fight to get the E/W petals up high enough to allow for the N/S petals to open.  This issue was likely caused by a failure with the compressor used to actuate the pistons used to cycle the petals.  The compressor tripped, likely leading to the petals closing in an uncontrolled order.  Compressor has been running reliably since the trip early yesterday morning.  That said, work has finally begun on plumbing in a “new” oilless compressor, purchased ~1 decade ago, that is housed out by the compressor used for the airbag system.  This will act as a replacement, moving the currently used compressor into the role of spare system.  This will provide added redundancy to the already in-place backup of N2 bottles.

As an aside, when working on getting the mirror covers open last night, I found an elastic rope connected on the far eastern edge of the southern petal.  Pulling on it broke the strap.  This appears to have no impact on system performance, so I am unclear as to what the purpose of the band was.  

Tested dome motion this morning and found that the encoder was not reading out CW motions.  This would lead to the symptoms described below.  My guess as to why the dome behaved for some targets after rehoming the system is that targets desired demanded dome motion in the CCW direction.  I cycled the connection to the dome encoder interface box, correcting the problem.  Numerous tests were run throughout the day to ensure that the system is functioning properly.